OpenStudy (iwanttogotostanford):

HEY NEED HELP ASAP PLEASE: Using the completing-the-square method, find the vertex of the function f(x) = –2x^2 + 12x + 5 and indicate whether it is a minimum or a maximum and at what point. Maximum at (–3, 5) Minimum at (–3, 5) Maximum at (3, 23) Minimum at (3, 23)

OpenStudy (zzr0ck3r):

\(f(x)=ax^2+bx+c\\f(x)=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}\) Has vertex \((-\frac{b}{2a}, c-\frac{b^2}{4a})\) This is a maximum if \(a<0\) and a minimum if \(a>0\)

OpenStudy (iwanttogotostanford):

i am stuck between A and B

OpenStudy (iwanttogotostanford):

@zzr0ck3r

OpenStudy (iwanttogotostanford):

@Nnesha

OpenStudy (iwanttogotostanford):

@Hero

OpenStudy (zzr0ck3r):

Do you not understand something I wrote?

OpenStudy (iwanttogotostanford):

NO i am just stuck between answer A and B

OpenStudy (zzr0ck3r):

Show me your work. i.e. show me your vertex form.

OpenStudy (iwanttogotostanford):

i just want to know if i got it right and i am stuck between A and B!

hero (hero):

For quadratic functions with the form f(x) = ax^2 +bx+c , if a is negative, then the function will have a maximum value at the vertex.

OpenStudy (zzr0ck3r):

that is not what we are here for.

OpenStudy (zzr0ck3r):

@hero must be on his phone, that took way to long and your just repeating :P

hero (hero):

Yeah, I don't have the best tools right now.

OpenStudy (iwanttogotostanford):

s o it is A?

hero (hero):

.Not tellin

OpenStudy (iwanttogotostanford):

i did the work but I'm just stuck between minimum and maximum which is A or B! p

hero (hero):

We've given you enough info.

OpenStudy (iwanttogotostanford):

i don't know yet.

OpenStudy (iwanttogotostanford):

how do i tell between minimum and maximum

hero (hero):

Read What I wrote about quadratic functions once more.

OpenStudy (iwanttogotostanford):

i did i still don't know A or B

OpenStudy (iwanttogotostanford):

@mathway

hero (hero):

This is similar to what happens when you give someone money, but they don't Know how to spend it.

OpenStudy (iwanttogotostanford):

who wouldn't know how to spend money

OpenStudy (iwanttogotostanford):

@mathway please help

OpenStudy (anonymous):

Have you found the vertex form?

OpenStudy (iwanttogotostanford):

well kind of, i am just stuck between a and B

OpenStudy (anonymous):

Can you show me the steps you did?

OpenStudy (iwanttogotostanford):

i have it all written on paper i can't take a picture..

OpenStudy (anonymous):

You can select the Draw button to write your work.

OpenStudy (iwanttogotostanford):

well for my last step i Got A!! finally figured how to do it:-)

OpenStudy (anonymous):

So you will not show your work then? Okay.

OpenStudy (iwanttogotostanford):

no, i got the answer- thanks anyways though! @mathway

OpenStudy (anonymous):

All right, but I won't say if you're right or wrong.

OpenStudy (iwanttogotostanford):

ok

OpenStudy (iwanttogotostanford):

i know I'm right so thank you though